# Non-Markovianity of subsystem dynamics in isolated quantum many-body systems

### v1

@@ Abstract
Results— In Fig.1, we show results for the evolution of trace distance between reduced density matrices of subsystems composed of one to four spins (multiple spins in contiguous blocks), separated in time by $\delta\!=\!1$ and $2$ seconds for a quench starting from the paramagnetic ground state at $(J,h_{x},h_{z})\!=\!(0.2,1,0)$ to the ferromagnetic side $(J,h_{x},h_{z})\!=\!(1,0.1,0.5)$. Note that the latter parameters are deep in the non-integrable regime. The highly non-monotonic behaviour of the trace distances in Fig.1 is a telltale signature of strong non-Markovianity in the dynamics of these subsystems. Moreover, the oscillatory behaviour in trace distances appear to be rather persistent. This, we believe, is also a signature of very slow relaxation and anomalous thermalization, since it stands to reason that strong non-Markovianity should in general be an obstruction to fast and efficient relaxation. Note also that at any fixed time $t$, the trace distances between the temporally-separated states of the larger subsystems are strictly higher than those of the smaller subsystems : this is simply due to the contractivity of any distance measures under the partial trace operation.
A more revealing picture may be obtained by defining a simple degree of non-Markovianity as the cumulative magnitude of the revivals (increases) of the trace distances (other definitions of such a degree are possible of course). Fig.2 shows this degree, the upper row of which reveals two facts : I) no discernible mathematical pattern is seen for the case of paramagnetic-to-ferromagnetic quench except for the fact that this degree assumes noticeably smaller values for the larger subsystems of three and four spins, signifying that the dynamics of smaller subsystems of one and two spins are considerably more non-Markovian and further larger subsystems have progressively lessening non-Markovianity in their dynamics, and II) a negligible degree of non-Markovianity is seen in the case of ferromagnetic-to-paramagnetic quench (about $50$ times smaller than the opposite case), although a more systematic dependence on the subsystem sizes is seen. The subsystems’ dynamics in the latter case may thus be considered as effectively Markovian. We have seen similar results for paramagnetic-to-paramagnetic and ferromagnetic-to-ferromagnetic quenches as well. It is quite surprising that these latter quenches apparently induce effectively Markovian dynamics of subsystems (at least as per the behaviour of distance measures between quantum states), since it seems more natural that subsystems of closed quantum many-body systems should naturally exhibit some notable degree of non-Markovianity in far from equilibrium times for any parameter regimes because the process of equilibration of a closed quantum many body system is a complex process with various subsystems interacting with (and acting as "environments" for) each other in a common pursuit to attain a globally equilibrated state and the dynamics of a given subsystem (especially one that is sufficiently small compared to the total system size) plus its complement can not in general be expected to flow monotonously (i.e., a monotonous flow of information out of the subsystem to its environment) to equilibration in general, even more so because the environments themselves are also evolving dynamically and the interaction strengths between intra-subsystem spins, intra-environment spins and inter-subsystem-environment spins are all equal to each other. Yet, it is the direction of quenches and the consequent restructuring and mixing of the quantum states that may be the most decisive factor on whether the ensuing dynamics of subsystems is Markovian or otherwise.
@@
Figure 3: First row - Non-monotonicities of TVD vs time for $\delta=1,2$ seconds for paramagnetic-to-ferromagnetic quench. Second Row - Degree of TVD non-monotonicities vs $\delta$ in paramagnetic-to-ferromagnetic quench (left), and ferromagnetic-to-paramagnetic quench (right). Third Row - Dependence of this degree on varying longitudinal field (left) and transverse field (right) for 4-RDMs in the former quench.
@@
Remarkably, a more organized behaviour in the corresponding quantities is shown by the classical probability distributions furnished by the eigenvalues of the RDMs. Let us hypothesize for the moment that under CPTP maps acting on a pair of quantum states (density matrices), the distance between the corresponding probability vectors furnished by their respective eigenvalues also shows a monotonically non-increasing behaviour [[Althoughintuitivelyappealing, wearenotawareofanysuchresultsestablishedinquantuminformationtheorytothebestofourknowledge.Inclassicalinformationtheory, distancemeasuresbetweenprobabilitydistributionsalsoshowmonotonicallynon-increasingbehaviourunderclassicalprocessesandconsequentinequalitiesareknownasdata-processinginequalities, seee.g.][, whosequantumcounterpartsarethemonotonicinequalitiesgoverningdistancesbetweenquantumstatesunderquantumchannels/CPTPmaps.Butwhatclassical"dataprocessing"aspectsarecontainedwithinquantumchannelsactingonquantumstatesisasubtlequestioninquantuminformationtheory.]Cover1999]. If this hypothesis is true, could it be that any violation of this monotonicity under the action of a family of CPTP maps (dynamical maps) is also a signature of some aspect of non-Markovian dynamics of the underlying quantum states? [[Itisintriguingtopositthatthequantumdynamicsofsubsystemsalsocarry"classical"aspectsinsuchamannerthattheeigenvaluesofrespectivereduceddensitymatricesmightshowtheclassically-expectedmonotonicitiesfordistancemeasuresbetweenthemiftheunderlyingquantumdynamicsisMarkovian, andaviolationofthosemonotonicitieswhentheunderlyingquantumdynamicsisnon-Markovianasperquantummeasures.Itistemptingtoregardthisasaclassicalcharacteristicofquantumnon-Markovianitybutinasensedifferentfrom][.Itwouldbeworthinvestigatingthissystematicallyinfuture.]Milz2020]. Be that as it may, if the above hypothesis is true, then a degree of violation of said monotonicity may be defined again by the cumulative magnitude of the revivals (increases) of the employed distance measure. Fig.3 shows the non-monotonicities of the total variation distances defined previously and the corresponding "degree of non-monotonicities" for the paramagnetic-to-ferromagnetic quench, which shows several remarkable facts. Firstly, in the second row, the degree for any considered subsystem is strongly oscillatory but decreasing on an average with increasing temporal separation $\delta$. Secondly, the degrees for ferromagnetic-to-paramagnetic quench (right, second row) are about $\sim 30$ times smaller than those for the paramagnetic-to-ferromagnetic quench (left, second row). Thirdly, for paramagnetic-to-ferromagnetic quench (left, second row), the $\delta$-values corresponding to the visible maxima and minima are nearly identical for all considered subsystems (notice however that the maxima corresponding to 1-RDMs are slightly shifted to the left relative to those corresponding to the other subsystems), and they appear to nearly coincide with each other already for subsystems as small as three and four spins. Note however that this degree at any given $\delta$ is larger in value for larger subsystems, in contrast with the degrees of non-Markovianity shown in Fig.2. Clearly then, this degree can not be a measure of non-Markovianity of quantum states in the usual sense. Fourthly, in the third row of Fig.3 a general increase of these degrees with increasing $h_{z}$ is seen (for most values of $\delta$), whereas a rather drastic reduction in their oscillatory amplitude is seen with increasing $h_{x}$. Once again, quite remarkably the maxima and minima of these degrees at varying values of $h_{z}$ and $h_{x}$ occur at nearby $\delta$-values, with relative shifts becoming more prominent at higher $\delta$-values (especially for the integrable case $h_{z}=0$). Moreover, there seem to exist remarkable timescales associated with the behaviour of TVD between the eigenvalues of considered RDMs : a timescale of $\sim 0.78(\pm 0.02)$ seconds in evolution times $t$ between the minima in the first row of Fig.3, which is largely independent of the subsystem size and of the temporal-separation $\delta$ (we checked this for several $\delta$ values), and a second timescale of $\sim\{1.55,1.6\}$ seconds in the temporal-separation $\delta$ between the maxima (or between the minima) emerges in the degrees of TVD non-monotonicities (second row) which also is largely independent of the subsystem size (additionally, even with varying transverse and longitudinal fields there is a timescale of approximately the same value but in a cruder fashion).
@@
We remarked earlier that the magnetically-ordered regime with non-zero longitudinal field is known to exhibit slow thermalization dynamics. This is mainly due to the existence of confining potential between certain excitations that separate domain walls [31, 32, 33, 34], as well as due to the presence of scars (non-thermalizing eigenstates) deep in the spectrum whose presence is related to and indeed enhanced by the presence of confinement [59, 41]. The strength of the confining potential is directly proportional to the longitudinal field and we have seen in this article that, while the trace distance degrees signifying subsystem non-Markovianity did not show sufficiently differential behaviour with increasing values of the longitudinal field, the degrees of non-monotonicities of total variation distance between eigenvalues of subsystem RDMs showed marked increase with increasingly higher longitudinal fields. This intriguing connection would be worth understanding more precisely in future. Additionally, the asymmetry between the two directions of quenches, seen in this work in the context of subsystem non-Markovianity, was also noted previously in the context of dynamical quantum phase transitions and differentiating dynamical behaviour was seen in quantum mutual information (between small subsystems) and in leading members of the entanglement spectrum [60].
@@
Concluding remarks— By numerically investigating the dynamical behaviour of trace distances between temporally-separated reduced density matrices corresponding to small subsystems of a closed quantum spin system in one dimension, and with a viewpoint of looking at these subsystems as open quantum systems, we have presented numerical evidence of non-Markovianity of these subsystems’ dynamics and the dependence of this characteristic on various factors at play. Notably, it was demonstrated that the direction (in parameter space) of quenching is a major factor for noticeable non-Markovianity, and that smaller subsystems showcase larger non-Markovianity in their dynamics. We have also revealed a remarkably systematic dynamical behaviour in the measure of distance between the corresponding eigenvalues of the considered reduced density matrices, the interpretation and explanation of which we can not offer at this time and will be left for future work. We hope this sort of a viewpoint can be put to practice more widely by applying some of the central tools and concepts of open quantum systems theory [22, 23] onto uncovering and understanding many facets of non-equilibrium dynamics of isolated quantum many-body systems. While we focused in this article on the Ising spin model for simplicity and the paradigmatic value held by this model, we have observed signatures of subsystem non-Markovianity in quench dynamics of Heisenberg spin systems also, albeit they were most pronounced in quenches from paramagnetic to magnetically ordered phases in those systems as well, hence our focus on the simpler Ising spin chain in this article. In a forthcoming article we will present results on the dynamical behaviour of other quantum informational quantities that accompany subsystem non-Markovianity, as well as an example of another simple spin system where, unlike the Ising model in this work, a quench from a magnetically ordered to a paramagnetic side (where the latter is topological in nature) also shows pronounced non-Markovianity. Further important questions that need to be addressed include the effects of performing measurements on subsystem non-Markovianity, whether effective spin-boson type models or collision model descriptions can be constructed to mimic the open system dynamics of subsystems of closed many-body systems, a characterization of the properties of the dynamical maps themselves, and whether other frameworks for characterizing and simulating non-Markovianity can provide additional or deeper insights. Work is in progress in some of these directions and we hope to be able to report on them in future.
@@
Acknowledgements— Financial support from the Department of Atomic Energy, India is gratefully acknowledged.

### v2

@@ IV Results and discussion for the mixed-field Ising spin chain
In Fig.1, we show results for the evolution of trace distance $T_{d}(\rho_{t+\delta}^{\ell},\rho_{t}^{\ell})$, for subsystems composed of one ($\ell\!=\!1$) to four spins in contiguous blocks ($\ell\!=\!4$), separated in time by $\delta\!=\!1$ and $2$ for a quench starting from the paramagnetic ground state at $(J,h_{x},h_{z})\!=\!(0.2,1,0)$ to the ferromagnetic side $(J,h_{x},h_{z})\!=\!(1,0.1,0.5)$ (in the labels of Fig.1, $k-$RDM refers to reduced density matrices (RDMs) of subsystem size, $k\!=\!\{1,2,3,4\}$). Note that the latter parameters are deep in the non-integrable regime. The highly non-monotonic behaviour of the trace distances in Fig.1 is a telltale signature of strong non-Markovianity in the dynamics of these subsystems. Moreover, the oscillatory behaviour in trace distances appear to be rather persistent. This, we believe, is also a signature of very slow relaxation and anomalous thermalization, since it stands to reason that strong non-Markovianity should in general be an obstruction to fast and efficient relaxation. Note also that at any fixed time $t$, the trace distances between the temporally-separated states of the larger subsystems are strictly higher than those of the smaller subsystems : this is simply due to the contractivity of any distance measures under the partial trace operation (a CPTP operation).
A more revealing picture may be obtained by plotting the degree of non-Markovianity Eq.(8) vs the temporal separation $\delta$. Fig.2 shows this degree, the upper row of which reveals two facts : I) no discernible mathematical pattern is seen for the case of paramagnetic-to-ferromagnetic quench except for the fact that this degree assumes noticeably smaller values for the larger subsystems of three and four spins, signifying that the dynamics of smaller subsystems of one and two spins are considerably more non-Markovian and further larger subsystems have progressively lessening non-Markovianity in their dynamics, and II) a negligible degree of non-Markovianity is seen in the case of ferromagnetic-to-paramagnetic quench (about $\sim 50$ (or more) times smaller than the opposite case), although a more systematic dependence on the subsystem sizes is seen. The subsystems’ dynamics in the latter case may thus be considered as effectively Markovian. We have seen similar results for paramagnetic-to-paramagnetic and ferromagnetic-to-ferromagnetic quenches as well.
@@
Figure 3: Strongly non-monotonic behaviour of total variation distances (TVD), Eq.(2) (between descendingly-ordered eigenvalues of $\delta-$separated RDMs of small subsystems), on the y-axes vs time for $\delta\!=\!1$ (left) and $\delta\!=\!2$ (right) in the paramagnetic-to-ferromagnetic quench $(J,h_{x},h_{z})\!=\!(0.2,1,0)\rightarrow(1,0.1,0.5)$. Here too, $k-$RDMs mean the same as in Fig.1. Note the existence of a timescale of $\sim 0.78(\pm 0.02)$ in evolution times $t$ between the minima, which is largely independent of the subsystem size and of the temporal-separation $\delta$.
Figure 4: Degree of TVD revivals $\mathcal{D}_{1}(\delta)$, Eq.(10), on the y-axes for various subsystems vs $\delta$. Upper Row - For various small subsystems in paramagnetic to ferromagnetic quench $(J,h_{x},h_{z})\!=\!(0.2,1,0)\rightarrow(1,0.1,0.5)$ (left), and ferromagnetic to paramagnetic quench $(J,h_{x},h_{z})\!=\!(1,0.1,0.5)\rightarrow(0.2,1,0)$ (right). Here too, $k-$RDMs mean the same as in Fig.1. Note that the degree of TVD revivals in the latter quench are roughly 30 times smaller than in the former. Particularly prominent in the the former quench is the existence of a timescale of $\sim\{1.55,1.6\}$ in the temporal-separation $\delta$ between the maxima (or between the minima) in the degree of TVD revivals, which also is largely independent of the subsystem size. Lower Row - Dependence of $\mathcal{D}_{1}(\delta)$ on varying longitudinal field $h_{z}$ (left) and transverse field $h_{x}$ (right) for 4-RDMs in the former quench.
Remarkably, a more organized behaviour in the corresponding quantities is shown by the classical probability distributions furnished by the eigenvalues of the RDMs. Let us hypothesize for the moment that under CPTP maps acting on a pair of quantum states (density matrices), the distance between the corresponding probability vectors furnished by their respective eigenvalues also shows a monotonically non-increasing behaviour. Although intuitively appealing, we are not aware of any such results established in quantum information theory to the best of our knowledge. In classical information theory, distance measures between probability distributions also show monotonically non-increasing behaviour under classical processes and consequent inequalities are known as data-processing inequalities, see e.g. [57], whose quantum counterparts are the monotonic inequalities governing distances between quantum states under quantum channels/CPTP maps. But what classical "data processing" aspects are contained within quantum channels acting on quantum states is a subtle question in quantum information theory. If this hypothesis is true, could it be that any violation of this monotonicity under the action of a family of CPTP maps (dynamical maps) is also a signature of some aspect of non-Markovian dynamics of the underlying quantum states? It is intriguing to posit (and would be worthwhile to investigate systematically in future) that the quantum dynamics of subsystems also carry "classical" aspects in such a manner that the eigenvalues of respective reduced density matrices might show the classically-expected monotonicities for distance measures between them if the underlying quantum dynamics is Markovian, and a violation of those monotonicities when the underlying quantum dynamics is non-Markovian as per quantum measures. It is tempting to regard this as a classical characteristic of quantum non-Markovianity but in a sense different from [58].
@@
Be that as it may, if the above hypothesis is true, then a degree of violation of said monotonicity may be defined again by the cumulative magnitude of the revivals (increases) of the employed distance measure. Fig.3 shows highly non-monotonic behaviour of the total variation distances $V_{d}(q_{t+\delta}^{\ell},q_{t}^{\ell})$ between descendingly-ordered sets of eigenvalues of RDMs (separated in time by $\delta\!=\!1,2$) corresponding to subsystems of size $\ell$, for the paramagnetic-to-ferromagnetic quench. In Fig.4, we show the corresponding degree of TVD revivals $\mathcal{D}_{1}(\delta)$ (Eq.10), which shows several remarkable facts : I) In the upper row, this degree for any considered subsystem is strongly oscillatory but decreasing on an average with increasing temporal separation $\delta$. II) The degrees for ferromagnetic-to-paramagnetic quench (right, upper row) are about $\sim 30$ times smaller than those for the paramagnetic-to-ferromagnetic quench (left, upper row). III) For paramagnetic-to-ferromagnetic quench (left, upper row), the $\delta$-values corresponding to the visible maxima and minima are nearly identical for all considered subsystems (notice however that the maxima corresponding to 1-RDMs are slightly shifted to the left relative to those corresponding to the other subsystems), and they appear to nearly coincide with each other already for subsystems as small as three and four spins. Note however that this degree at any given $\delta$ is larger in value for larger subsystems, in contrast to the degrees of non-Markovianity shown in Fig.2. Clearly then, this degree can not be a measure of non-Markovianity of quantum states in the usual sense. IV) In the lower row of Fig.4 a general increase of these degrees with increasing $h_{z}$ is seen (for most values of $\delta$), whereas a rather drastic reduction in their oscillatory amplitude is seen with increasing $h_{x}$. Once again, quite remarkably the maxima and minima of these degrees at varying values of $h_{z}$ and $h_{x}$ occur at nearby $\delta$-values, with relative shifts becoming more prominent at higher $\delta$-values (especially for the integrable case $h_{z}=0$).
Moreover, there seem to exist remarkable timescales associated with the behaviour of TVD between the eigenvalues of considered RDMs : a timescale of $\sim 0.78(\pm 0.02)$ in evolution times $t$ between the minima in Fig.3, which is largely independent of the subsystem size and of the temporal-separation $\delta$ (we checked this for several $\delta$ values), and a second timescale of $\sim\{1.55,1.6\}$ in the temporal-separation $\delta$ between the maxima (or between the minima) emerges in the degrees of TVD revivals in Fig.4 (upper row), which also is largely independent of the subsystem size (additionally, even with varying transverse and longitudinal fields there is a timescale of approximately the same value but in a cruder fashion, as seen in lower row of Fig.4).
V More discussion
@@
We remarked earlier that the magnetically-ordered regime with non-zero longitudinal field is known to exhibit slow thermalization dynamics. This is mainly due to the existence of confining potential between certain excitations that separate domain walls [37, 38, 39, 40], as well as due to the presence of scars (non-thermalizing eigenstates) deep in the spectrum whose presence is related to and indeed enhanced by the presence of confinement [59, 46]. The strength of the confining potential is directly proportional to the longitudinal field and we have seen in this article that, while the trace distance degrees signifying subsystem non-Markovianity did not show sufficiently differential behaviour with increasing values of the longitudinal field, the degrees of non-monotonicities of total variation distance between eigenvalues of subsystem RDMs showed marked increase with increasingly higher longitudinal fields. This intriguing connection would be worth understanding more precisely in future. Additionally, the asymmetry between the two directions of quenches, seen in this work in the context of subsystem non-Markovianity, was also noted previously in the context of dynamical quantum phase transitions and differentiating dynamical behaviour was seen in quantum mutual information (between small subsystems) and in leading members of the entanglement spectrum [60], and many more quantum informationally distinguishing differences between the quench directions have been reported in [61]. Both of these works have highlighted the driving role played by entanglement in the pronounced features shown in the paramagnetic-to-ferromagnetic quenches as opposed to the opposite quench.
Despite the weak nearest-neighbor couplings in the paramagnetic side already earmarking subsystem dynamics to be within the Born-Markov paradigm, but noting that the interaction strengths between intra-subsystem spins, intra-environment spins and inter-subsystem-environment spins are also all equal (or comparable) to each other, with the environment also simultaneously being far from equilibrium, it may appear somewhat surprising that the ferromagnetic-to-paramagnetic quenches induce effectively Markovian dynamics of subsystems (at least as per the behaviour of distance measures between quantum states) and at least marginally noticeable (even if weakened) non-Markovianity may have been expected a priori, since it seems more obvious to expect that subsystems of closed quantum many-body systems should naturally exhibit some notable degree of non-Markovianity in the far-from-equilibrium times for quenches between any parameter regimes because the process of equilibration of a closed quantum many body system is a complex process, with various subsystems interacting with (and acting as "environments" for) each other in a common pursuit to attain a globally equilibrated state. Therefore one might argue that the dynamics of a given subsystem (especially one that is sufficiently small compared to the total system size) plus its complement can not in general be expected to flow monotonously (i.e., a monotonous flow of information out of the subsystem to its environment) to equilibration in general, even more so because the environments themselves are also evolving dynamically. Yet, as we have seen it is the direction of quenches and the consequent restructuring and mixing of the quantum states that may be the most decisive factor on whether the ensuing dynamics of subsystems is Markovian or otherwise (in fact, general arguments suggest that subsystem dynamics being effectively Markovian should be more typical in most quantum systems’ closed dynamics [20]). These are crucially affected by the underlying physical scenarios as discussed in the previous paragraphs, such as the nature of entanglement creation needed to approach the target state, and the freeness or confinement of excitations to propagate between the subsystems in order to bring about the needed mixing and restructuring of the Hilbert space. Thus, the quasiparticle picture of quench dynamics [62, 63, 64], where the elementary excitations propagate almost ballistically to bring about a linear growth of entanglement and fast equilibration dynamics, is the probable cause of the effectively Markovian dynamics of subsystems in the ferromagnetic-to-paramagnetic quenches. It is already known that this picture is violated in models with confinement, thus linking satisfaction or violation of this picture with consequent Markovianity or its lack thereof of the dynamics of subsystems of closed quantum many-body systems.
Furthermore, whenever subsystem dynamics is essentially Markovian, it is in principle plausible and would be worthwhile to derive Markovian master equation based descriptions of subsystems’ dynamics and consequently develop an understanding of equilibration or thermalization and the ETH picture from this perspective. Recently, thermalization and ETH in open quantum systems have been studied in several contexts [65, 66, 67, 68, 69], and it is hoped that similar descriptions could be extended to the understanding of approach to thermalization and ETH in closed quantum many-body systems from this perspective. A different approach to a subsystem based understanding of equilibration, thermalization and the emergence of statistical mechanics in closed quantum many-body systems is already developed in e.g. [70, 71, 72, 73, 74].
VI Conclusion
@@
By numerically investigating the dynamical behaviour of trace distances between temporally-separated reduced density matrices corresponding to small subsystems of a closed quantum spin system in one dimension, and with a viewpoint of looking at these subsystems as open quantum systems, we have presented numerical evidence of non-Markovianity of these subsystems’ dynamics and the dependence of this characteristic on various factors at play. Notably, it was demonstrated that the direction (in parameter space) of quenching is a major factor for noticeable non-Markovianity, and that smaller subsystems showcase larger non-Markovianity in their dynamics in this quench dynamics (contrary to the results of [20] that are more generally applicable, to which it seems the "extreme" paramagnetic-to-ferromagnetic quenches apparently constitute a (plausibly small) set of exceptions). We have also revealed a remarkably systematic dynamical behaviour in the measure of distance between the corresponding eigenvalues of the considered reduced density matrices, the interpretation and explanation of which we can not offer at this time and will be left for future work. We hope this sort of a viewpoint can be put to practice more widely by applying some of the central tools and concepts of open quantum systems theory [28, 29] onto uncovering and understanding many facets of non-equilibrium dynamics of isolated quantum many-body systems (see also a recent work in this spirit [75]). Specific to the issue of subsystem non-Markovianity described in this work, it would be worthwhile in future to derive an analytical description of these results perhaps along the lines of e.g. [76, 77, 78, 79] by invoking the properties of generalized Bloch balls [80].
While we focused in this article on the Ising spin model for simplicity and the paradigmatic value held by this model, we have observed similar signatures of subsystem non-Markovianity in quench dynamics of Heisenberg spin systems also, albeit they were most pronounced in extreme quenches from the paramagnetic to magnetically ordered phases in those systems as well, hence our focus on the simpler Ising spin chain in this article. In a recent work [61], we have revealed certain fine-grained attributes of entanglement dynamics for the same quench dynamics as in this work, and these go (possibly necessarily) hand-in-hand with subsystem non-Markovianity. Further important questions that need to be addressed include the effects of performing measurements on subsystem non-Markovianity, whether effective spin-boson type models or collision model descriptions can be constructed to mimic the open system dynamics of subsystems of closed many-body systems, a characterization of the properties of the dynamical maps themselves, and whether other approaches for characterizing and simulating non-Markovianity can provide additional or deeper insights. Work is in progress in some of these directions and we hope to be able to report on them in future.
@@
Acknowledgements— We thank the Department of Atomic Energy, India for financial support.